# An effective Bombieri–Vinogradov error term for sifting problems

### v1

@@ 4.1  Preliminary lemmas
so that the condition $|\operatorname{\mathcal{A}}_{d}|\sim g(d)|\operatorname{\mathcal{A}}|$. More generally, the Chinese remainder theorem gives that the condition (4.1) holds with $\gamma_{1}=2$. We then have for $z_{2}>z_{1}>3$},
@@
Applying Mertens’ first theorem to (5.2) then gives
@@
$R_{DHR}(D)=2\sum_{\begin{subarray}{c}d<D\\ d\mid P(z)\end{subarray}}4^{\omega(d)}|r(d)|\leq\sum_{\begin{subarray}{c}d<D\\ d\mid P(z)\end{subarray}}8^{\omega(d)}|r(d)|,$ (5.12)
@@
Therefore, with $X=|\operatorname{\mathcal{A}}|\sim\pi(\sqrt{n})$ and $V(z)$ bounded as in (5.10), we can apply Theorem 1.7 with $A=2$, $\operatorname{\mathcal{X}}=\sqrt{n}$, $B=265$, $\gamma=16$ and $s=4.267$ to yield an effective bound
@@
If one were to make the above proof explicit, then the value of $N$ for which $S(\operatorname{\mathcal{A}},\operatorname{\mathcal{P}},z)>0$ for all $n>N$, would be very large. This is because of the value of $B=266$ which would require one to use a very good bound for the error term in the prime number theorem for arithmetic progressions (see Lemma 2.5). To be able to use a lower value of $B$, one should use a sieve that requires as small a value of $\gamma$ as possible. For example, if the 2-dimensional Rosser-Iwaniec sieve [13, Theorem 1] was used instead of (5.11), then one could take $\gamma=2$ and $B=14$. However, this would come at the cost of slightly increasing the number of prime factors of $n-q^{2}$.
@@ Acknowledgements
The author thanks Bryce Kerr for his comments on a preliminary version of this paper.

### v2

@@ 4.1  Preliminary lemmas
so that $|\operatorname{\mathcal{A}}_{d}|\sim g(d)|\operatorname{\mathcal{A}}|$. More generally, the Chinese remainder theorem gives that the condition (1.21) holds with $\gamma_{1}=2$. We then have for $z_{2}>z_{1}>3$},
@@
Applying Mertens’ theorem to (5.2) then gives
@@
$R_{DHR}(D)=2\sum_{\begin{subarray}{c}d<D\\ (d,\overline{P})=1\end{subarray}}4^{\omega(d)}|r(d)|\leq\sum_{\begin{subarray}{c}d<D\\ (d,\overline{P})=1\end{subarray}}8^{\omega(d)}|r(d)|,$ (5.12)
@@
Therefore, with $X=|\operatorname{\mathcal{A}}|\sim\pi(\sqrt{n})$ and $V(z)$ bounded as in (5.10), we can apply Theorem 1.7 with $A=2$, $\operatorname{\mathcal{X}}=\sqrt{n}$, $B=266$, $\gamma=16$ and $s=4.267$ to yield an effective bound
@@
If one were to make the above proof explicit, then the value of $N$ for which $S(\operatorname{\mathcal{A}},\operatorname{\mathcal{P}},z)>0$ for all $n>N$ would be very large. This is because of the value of $B=266$ which would require one to use a very good bound for the error term in the prime number theorem for arithmetic progressions (see Lemma 2.5). To be able to use a lower value of $B$, one should use a sieve that requires as small a value of $\gamma$ as possible. For example, if the 2-dimensional Rosser–Iwaniec sieve [15, Theorem 1] was used instead of (5.11), then one could take $\gamma=2$ and $B=14$. However, this would come at the cost of slightly increasing the number of prime factors of $n-q^{2}$.
@@ Acknowledgements
The author thanks the anonymous referees for their useful comments. The author also thanks Bryce Kerr for his comments on a preliminary version of this paper.